Solve the formula for the specified variable.
V = 1/3Bh for h.
Answer choices:
A. h = V/3B
B. h = B/3V
C. h = 3B/V
D. h = 3V/B

Answer by stanbon(75887)

V = 1/3Bh for h
3V = Bh
h = (3V)/B
---
Ans: D
=============
cheers,
stan H.
==============
Answer choices:
A. h = V/3B
B. h = B/3V
C. h = 3B/V
D. h = 3V/B